40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 16 17:02:41 2013, Alex Cole, Configuration, Electronics, AS Table Additions photo_1_(1).JPGphoto_2_(2).JPG
    Reply  Wed Jul 17 11:13:36 2013, Alex Cole, Configuration, Electronics, AS Table Additions photo_(1).JPG
       Reply  Wed Jul 17 16:15:42 2013, Alex Cole, Configuration, Electronics, AS Table Additions photo_1_(3).JPGphoto_2_(3).JPGphoto_3.JPG
Message ID: 8863     Entry time: Wed Jul 17 16:15:42 2013     In reply to: 8862
Author: Alex Cole 
Type: Configuration 
Category: Electronics 
Subject: AS Table Additions 

[Eric, Alex]

We decided that the POY Table would be a better home for our REF DET (Newport 1611 FC-AC) than the AS Table. We moved the PD to the POY Table (1st attachment) and routed a fiber from our 1x16 Optical Splitter in the OMC_North rack to the POY Table. REF DET's power supply is now located under the POY table (2nd attachment). We left the fiber described in the previous post on the AS Table.

Afterwards, we hooked a fiber up to our laser module to test it (3rd attachment). The laser was not being distributed, just going to one fiber with a power meter at its end. Everything turns out, but we realized we need to read the power supply's manual before continuing. 

 

 

Attachment 1: photo_1_(3).JPG  113 kB  | Hide | Hide all
photo_1_(3).JPG
Attachment 2: photo_2_(3).JPG  123 kB  | Hide | Hide all
photo_2_(3).JPG
Attachment 3: photo_3.JPG  123 kB  | Hide | Hide all
photo_3.JPG
ELOG V3.1.3-